OpenStudy (loser66):

what is the condition of a to get \[ax_1\bar x_2+\bar a \bar x_1x_2>0\] where \(a,x_1,x_2 \in C\) \(x_1\neq 0\) \(x_2\neq 0\) Please, help

OpenStudy (loser66):

@phi

OpenStudy (phi):

I looked at this, and did not make much progress. If you ever find out, please post the answer here and tag me.

OpenStudy (loser66):

:( and :)

OpenStudy (loser66):

@satellite73

OpenStudy (p0sitr0n):

x bar is the mean?

OpenStudy (loser66):

conjugate of x

OpenStudy (anonymous):

Even if both \(x\) values are non-zero it is still possible that they multiply to 0, isn't it?

OpenStudy (loser66):

not sure.

OpenStudy (anonymous):

What about \((1-i)(-1+i)\)

OpenStudy (anonymous):

Oh wait, I guess that one doesn't matter.

OpenStudy (anonymous):

Let me think... \[ (a+bi)(c+di) = (ac-bd) + (ad+bc)i \]So we have: \[ ac=bd\\ ad=-bc \]This means \(a = bd/c\). so \[ bd^2/c =bc \implies bd^2=-bc^2\implies d^2=-c^2 \]The square of real numbers are positive so it really is impossible.

OpenStudy (anonymous):

I have a question....

OpenStudy (anonymous):

How can imaginary number be greater than another number?

OpenStudy (anonymous):

Is \(i > 0\)?

OpenStudy (loser66):

i cannot >0

OpenStudy (anonymous):

So this product has to be a real number...

OpenStudy (anonymous):

actually!! That makes sense. You are adding two conjugates. They must add up to real number.

OpenStudy (anonymous):

Maybe we can simplify the problem to: \[ ax+\bar{a}\bar{x}>0 \]

OpenStudy (anonymous):

It must be the case that \(2\Re(ax) > 0 \implies \Re(ax) > 0 \)

OpenStudy (anonymous):

This means \[ \Re(a)\Im(x)-\Re(x)\Im(a) > 0 \]

OpenStudy (loser66):

we have 3 numbers, not 2.

OpenStudy (anonymous):

I am saying let \(x=x_1\bar{x_2}\)

OpenStudy (loser66):

Can you give me a particular example to make it clear? Please

OpenStudy (anonymous):

Hmmmm, um.

OpenStudy (anonymous):

I don't think it can be simplified.